OpenStudy (anonymous):

write an equation for the line that is perpendicular to the given line and that passes through the given point -10x+8y=3;(15,12)

OpenStudy (anonymous):

Start by finding the slope of the line.

OpenStudy (anonymous):

How do I do that

OpenStudy (anonymous):

-10x+8y=3....slope is 5/4. You are looking for a line that is perpendicular to that slope, so that means the new slope would be -4/5

OpenStudy (anonymous):

now use that new slope and the given point (15,12) and plug it into the point slope formula.

OpenStudy (anonymous):

so, (y-12)=-4/5(x-15)

OpenStudy (anonymous):

so the new equation would be:

OpenStudy (anonymous):

y=-4/5x+24

OpenStudy (anonymous):

How did u get the slope

OpenStudy (anonymous):

Hello mr how do you find the slope of the line :)

OpenStudy (anonymous):

solve the equation for y. -10x+8y=3 -10x+8y=3 add 10x to both sides 8y=10x+3

OpenStudy (anonymous):

now divide both sides by 8

OpenStudy (anonymous):

y=(10x+3)/8 or 10/8x+3/8 this reduces to y=5/4x+3/8

OpenStudy (anonymous):

This new equation: y=5/4x+3/8 is in y intercept form, where y=mx+b, m represents the slope.....5/4

OpenStudy (anonymous):

Yay I got that. Sorry I would have u a medal but I can't now when I am gonna be on the desk top then I'll be your fan ~thanks so much :)

OpenStudy (anonymous):

Is that it ?

OpenStudy (anonymous):

Hello???

OpenStudy (anonymous):

Okay so now I'm not ont desktop but I just realized I can give u amedal with the phone